OpenStudy (anonymous):

Part 1: Create your own factorable trinomial. (1 point) Part 2: Explain, in complete sentences, how the trinomial is factored. (1 point) Part 3: Explain, in complete sentences, the process used to check the factors for accuracy. (1 point)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

I came up with the equation part, but I can't remember how to factor it... The equation is x^2+2x+1

jimthompson5910 (jim_thompson5910):

x^2+2x+1 factors to (x+1)(x+1) since 1+1 = 2 and 1*1 = 1

jimthompson5910 (jim_thompson5910):

another example x^2 + 5x + 6 factors to (x+3)(x+2) since 3+2 = 5 and 3*2 = 6

OpenStudy (anonymous):

Thank you soooo much!!!

jimthompson5910 (jim_thompson5910):

yw